Korup National Park (01.01.2014 to 03.31.2014)

You are here

PDF versionPDF version
Year: 
2014
Quarter: 
January - March
New Species: 

The African civet (Civettictis civetta) was photographed by a camera trap in array 1. It had not been seen on the camera traps in the previous 2 years. It is likely that the individual on the image is one of two that came from the Limbe Wildlife Center and were released on 28 December 2013, near the Ekundu Kundu trail, around Kilometer 3. This is not far from the area where this species was captured on image.

 

Protocol Activities: 

Climate station:

Climate data for the January – March 2014 were collected and uploaded to the TEAM website. The update of the climate station operating system and mounting of the third RH/T sensor were done Wednesday 9 April 2014

 

Trees and Lianas

The Third recensus of the six 1-ha vegetation plots was carried out in December 2013 and data entry is on-going.

 

Terrestrial vertebrates

Camera traps were successfully deployed in the two TEAM arrays in the southern part of the Korup National Park between January and March. Importation of the pictures into the deskTEAM and their annotation is ongoing. A new hunting protocol that CTFS has developed was implemented in the Park to quantify the density of hunting signs

 

 

Protocol Problems: 

On Array 1, only 28 of the 30 Camera traps deployed functioned properly during the 30 days. One camera trap failed to take any image, including the set up photos, while it did not indicate any problem during the setup. The second functioned for only two weeks and was indicating low batteries when picked up.  

In Array 2, a camera was damaged by a hunter. The camera was opened and water entered inside which caused the battery fittings to start rusting.

Some of the camera traps delivered blurry colour images which seem to be caused by humidity inside the cameras. Even though careful measures are taken to protect the cameras from humidity they display some problems with the digital screen and or blurry colour images (infrared images are fine).

 

Logistical Problems: 

The damage to the climate station - the cable connecting the rainfall gauge was cut - was repaired.

We had problems uploading the array 1 tpk files to the TEAM website from Cameroon, even with a fast internet connection. It was suggested that the data be sent to the US in a flash drive.

Suggestions: 

Guidelines are needed for the prevention of camera failure due to humidity.